# AppellateGen: A Benchmark for Appellate Legal Judgment Generation

### v2

@@ Abstract
AppellateGen: A Benchmark for Appellate Legal Judgment Generation
Hongkun Yang1,2thanks: Equal contribution.thanks: Work done while Hongkun Yang was an intern at Nanyang Technological University., Lionel Z. Wang1,311footnotemark: 1thanks: Project Leader., Wei Fan4, Yiran Hu5, Lixu Wang1, Chenyu Liu3, Shenghong Fu3, Haoyang Li3, Xin Xu3, Jiexin Zheng6, Wei Dong1thanks: Corresponding author. 1Nanyang Technological University, 2Ocean University of China, 3The Hong Kong Polytechnic University, 4Hong Kong University of Science and Technology, 5Tsinghua University, 6Peking University
@@ 1 Introduction
This distinction is particularly pronounced in Civil Law jurisdictions. Here, appellate courts operate under the principle of continuation of trial, allowing for the re-examination of facts and the admission of new evidence (apple1995primer). This procedural feature necessitates models capable of processing dynamic factual updates. According to judicial statistics from the Supreme People’s Court of China, the reversal rate of second-instance cases consistently hovered around 10% to 11% (exceeding 637,000 cases) from 2021 to 2023 SPC2021; SPC2022; SPC2023. This substantial volume of modified verdicts underscores the complexity inherent in appellate adjudication (luneburg1981specially; bruhl2010deciding), which demands complex reasoning to identify discrepancies and produce a binding judgment that affirms, reverses, or remands the initial ruling.
@@
Despite the importance of appellate review, existing legal benchmarks exhibit limitations in Task Formulation and Contextual Granularity. First, regarding formulation, existing benchmarks primarily assess discriminative capabilities through multiple-choice questions or classification chalkidis-etal-2022-lexglue; fei-etal-2024-lawbench. While recent initiatives have integrated generative tasks li-etal-2025-legalagentbench; li2024lexeval, they predominantly focus on drafting documents from static facts, overlooking the dialectical nature of appeals. Second, in terms of granularity, current methodologies typically model legal documents as isolated data points. Retrieval-centric benchmarks pipitone2024legalbench and document generation datasets li2025casegen generally treat case processing as a single-step inference task. This approach neglects the procedural interdependency between adjudicatory stages, specifically the dependency between initial judgments and appellate reviews.
To bridge this gap, we introduce AppellateGen, the first benchmark dedicated to Second-Instance Legal Judgment Generation. Unlike previous benchmarks, AppellateGen models the complete dispute lifecycle. We curate 7,351 paired cases, linking first-instance judgments with their corresponding appellate outcomes. We enhance this with a fine-grained annotation schema that identifies points of contention, extracts new facts introduced during the appeal, and categorizes the rationale for reversal.
@@
To validate the challenge posed by AppellateGen and investigate robust solutions, we propose the judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS) as a strong baseline. We observe that standard LLMs often struggle when processing the conflicting narratives and long-context evidence inherent in appellate review, leading to hallucinations. To address this, SLMAS imposes a Standard Operating Procedure (SOP) inspired by judicial workflows, decomposing the generation task into discrete stages: Analysis, Search, Predict, and Write. By orchestrating agents specialized in issue identification, retrieval, prediction, and drafting, our framework explicitly models the intermediate reasoning states required for valid adjudication. Empirical results demonstrate that this modular approach significantly enhances logical consistency and reversal prediction accuracy, surpassing both commercial and domain-specific baselines.
@@ 2 Related Work
2.1 Legal Judgment Prediction and Case Document Generation
Legal Judgment Prediction (LJP) has progressed from utilizing hierarchical legal structures (luo2017learning) and few-shot learning (hu2018few) to deploying attention-based architectures (chen2019charge; li2019mann; kang2019creating) and LLMs (nigam2024rethinking) for outcome deduction. Despite these advancements, the field has predominantly prioritized prediction over full document generation. While recent initiatives like CaseGen (li2025casegen) have pivoted toward generation, they remain restricted to first-instance trials, neglecting the appellate stage. Within the Chinese judicial framework, second-instance judgments are pivotal for their supervisory role, correcting errors from lower courts and unifying adjudication standards. To address this gap, we introduce the task of Second-instance Legal Judgment Generation.

### v3

@@ 1 Introduction
This distinction is pronounced in Civil Law jurisdictions. Here, appellate courts operate under the principle of continuation of trial, allowing for the re-examination of facts and the admission of new evidence (apple1995primer). This procedural feature necessitates models capable of processing dynamic factual updates. According to judicial statistics from the Supreme People’s Court of China, the reversal rate of second-instance cases consistently hovered around 10% to 11% (exceeding 637,000 cases) from 2021 to 2023 (SPC2021; SPC2022; SPC2023). This substantial volume of modified verdicts underscores the complexity inherent in appellate adjudication (luneburg1981specially; bruhl2010deciding), which demands complex reasoning to identify discrepancies and produce a binding judgment that affirms, reverses, or remands the initial ruling.
@@
Despite the importance of appellate review, existing legal benchmarks exhibit limitations in Task Formulation and Contextual Granularity. First, regarding formulation, existing benchmarks primarily assess discriminative capabilities through multiple-choice questions or classification (chalkidis-etal-2022-lexglue; fei-etal-2024-lawbench). While recent initiatives have integrated generative tasks (li-etal-2025-legalagentbench; li2024lexeval), they predominantly focus on drafting documents from static facts, overlooking the dialectical nature of appeals. Second, in terms of granularity, current methodologies typically model legal documents as isolated data points. Retrieval-centric benchmarks (pipitone2024legalbench) and document generation datasets (li2025casegen) generally treat case processing as a single-step inference task. This approach neglects the procedural interdependency between adjudicatory stages, specifically the dependency between initial judgments and appellate reviews.
@@
To bridge this gap, we introduce AppellateGen, the first benchmark dedicated to Second-Instance Legal Judgment Generation. Unlike previous benchmarks, AppellateGen models the complete dispute lifecycle. We curate 7,351 paired cases, linking first-instance judgments with their corresponding appellate outcomes. We enhance this with a fine-grained annotation schema that identifies points of contention, extracts new facts introduced during the appeal, and categorizes the rationale for reversal. To validate the challenge posed by AppellateGen and investigate robust solutions, we propose the judicial Standard Operating Procedure (SOP)-based Legal Multi-Agent System (SLMAS) as a strong baseline. We observe that standard LLMs often struggle when processing the conflicting narratives and long-context evidence inherent in appellate review, leading to hallucinations. To address this, SLMAS imposes a Standard Operating Procedure (SOP) inspired by judicial workflows, decomposing the generation task into discrete stages: Analysis, Search, Predict, and Write. By orchestrating agents specialized in issue identification, retrieval, prediction, and drafting, our framework explicitly models the intermediate reasoning states required for valid adjudication. Empirical results demonstrate that this modular approach significantly enhances logical consistency and reversal prediction accuracy, surpassing both commercial and domain-specific baselines.
@@ 2 Related Work
2.1 From Legal Prediction to Appellate Generation via Multi-Agents
Legal Judgment Prediction (LJP) has progressed significantly, evolving from hierarchical structures (luo2017learning) and few-shot learning (hu2018few) to deploying attention-based architectures (chen2019charge; li2019mann; kang2019creating) and LLMs (nigam2024rethinking). Recently, the field has witnessed two major paradigm shifts to address complex legal tasks. First, there is a pivot from mere outcome prediction to full document generation, exemplified by initiatives like CaseGen (li2025casegen). Second, to address single-agent limitations such as hallucinations, Multi-Agent Systems (MAS) have been introduced through collaborative simulation (ji2023towards; li2023camel). Frameworks like ChatLaw (cui2023chatlaw) and logic-graph-enhanced models (yuan2026multi) now leverage such role-playing collaboration to optimize judgment prediction and legal consultation. Despite these advancements, existing generative models and multi-agent frameworks share a critical limitation: they are inherently designed for first-instance trials. To bridge this gap, we introduce the task of Second-instance Legal Judgment Generation.